OpenStudy (anonymous):

Simplify −(6p + 12) 6p − 12 6p + 12 −6p + 12 −6p − 12

OpenStudy (anonymous):

i got the answer : -6p + -12 Am i right?

OpenStudy (anonymous):

i dont see it up there^^^^

OpenStudy (ajprincess):

\((+)*(-)=(-)\) -(6p+12)=-1(6p+12) =-1*6p+(-1)*12 =-6p-12 Is that clear?

mathslover (mathslover):

-(a+b) = \(\large{-a-b}\)

OpenStudy (anonymous):

but isint it when u multiply or divide 2 numbers with a different sign then the answer is negative?

mathslover (mathslover):

note that this is distributive property since we have: \[\large{-1(a+b)=-1(a)-1(b)=-1a-1b = -a-b}\]

OpenStudy (anonymous):

i dont get it ^^

mathslover (mathslover):

what is 6(a+b) ?

OpenStudy (anonymous):

wat i got was -6p + -12

OpenStudy (ajprincess):

When u multiply +12 by -1 it becomes -12. -(6p+12) =-1(6p+12) =(-1)*(+6p)+(-1)*(+12) =-6p-12 Is it clear nw?

OpenStudy (anonymous):

o ok its clear but the only thing that threw me off was that there was no plus sign between -6p and -12

OpenStudy (ajprincess):

there cannot be two signs within two terms.

OpenStudy (anonymous):

wat do u mean by that? im sorry i keep asking questions i just wanna make sure i understand wat im learning so i dont have problems in the future

OpenStudy (ajprincess):

\(a+-b\) is wrong. \(a+b\) or \(a-b\) is right.

OpenStudy (anonymous):

ok but wouldnt a person be confused if there was -6p -12 and think it was subtraction wen it should have been addition?

OpenStudy (anonymous):

if there was -6p-12 then it is subtraction.

OpenStudy (anonymous):

you are right when you Multiply or divide two no with different sign answer is negative thats why when -1 is Multiplied with +12 then it comes out as -12.

OpenStudy (ajprincess):